# Multi-Atlas Brain Network Classification through Consistency Distillation and Complementary Information Fusion

### v1

@@ 4.3 Subject- and Population-level Consistency
Once we obtain the cluster assignment matrix $\boldsymbol{S}$ and the embedded node feature matrix $\boldsymbol{Z}$, we generate a new feature matrix $\hat{\boldsymbol{H}}\in\mathbb{R}^{n^{\prime}\times d}$ by $\hat{\boldsymbol{H}}=\boldsymbol{S}^{\mathsf{T}}\boldsymbol{Z}$. This coarsening process can reduce the number of nodes to get higher-level node representations. To avoid GNN treating each ROI and each node cluster equally, we adopt an entropy regularization to the assignment matrices of each atlas:
$\small\mathcal{L}_{E}=\frac{1}{n^{\prime}}\sum^{n^{\prime}}_{i=1}\left(\operatorname{entropy}(\boldsymbol{S}^{a}[i,:])+\operatorname{entropy}(\boldsymbol{S}^{b}[i,:])\right),\operatorname{entropy}(\boldsymbol{p})=-\sum_{j=1}^{n^{\prime}}\boldsymbol{p}_{j}\log(\boldsymbol{p}_{j}).$ (9)
We elaborate on the module of subject-level consistency in Figure 1d. Through two DiffPool layers, we produce high-quality representations for each atlas by extracting high-level node representations. Then we are able to apply a contrastive loss to them by considering representations from the same subject as positive pairs $\mathcal{P}^{pos}=\{(\hat{\boldsymbol{H}}^{a}_{i},\hat{\boldsymbol{H}}^{b}_{i})\}:i=1,\dots,bz\}$ and representations from different subjects as negative pairs $\mathcal{P}^{neg}=\{(\hat{\boldsymbol{H}}^{a}_{i},\hat{\boldsymbol{H}}^{b}_{\neg i}):i=1,\dots,bz\}$:
@@
Population-level Consistency. The readout function $\boldsymbol{m}=\operatorname{READOUT}\left(\boldsymbol{H}\right)$ is an essential component of learning the graph-level representations $\boldsymbol{m}\in\mathbb{R}^{d}$ for brain network analysis (e.g., classification), which maps a set of learned node-level embeddings to a graph-level embedding. To further constrain the consistency for graph representations across different atlases, we introduce a mean squared error (MSE) loss on the population level. As shown in Figure 1e, a population graph $\boldsymbol{G}$ is constructed by computing the similarity of each two subjects’ graph representations in the same atlas. The intuition here is we aim to maintain the relationship of subjects across atlases, instead of directly enforcing graph representations of two atlases to be the same. Such loss is formulated as follows:
@@
Total Loss. The model is supervised by a commonly-used cross-entropy loss $\mathcal{L}_{cls}$ [9] for graph classification. The total loss is computed by:
@@
5 Experimental Results
5.1 Brain Network Datasets
Table 1: Statistics of Brain Network Datasets.
Dataset Condition Subject# Class# Class Name
ABIDE Autism Spectrum Disorder 1025 2 {TC, ASD}
ADNI Alzheimer’s Disease 1326 6 {CN, SMC, MCI, EMCI, LMCI, AD}
PPMI Parkinson’s Disease 209 4 {NC, SWEDD, prodromal, PD}
Matai Mild Traumatic Brain Injury (mTBI) 60 2 {pre-season, post-season}
We use four brain network datasets from different data sources for various disorders, which are ABIDE [10] for Autism (ASD), ADNI [11] for AD, PPMI [3] for Parkinson’s disease (PD), and Mātai for mild traumatic brain injury (mTBI) [54]. Statistics of the brain network datasets are summarized in Table 1. The atlases we use are Schaefer [44] and AAL [47] with 100 and 116 ROIs, respectively. The detailed dataset description is provided in Appendix D while the implementation detail of our experiments is given in Appendix E.
5.2 Baseline Models
@@
We use 7 single-atlas methods and 6 multi-atlas methods as baselines to evaluate our proposed AIDFusion, including: (1) Conventional machine learning (ML) models: Logistic Regression (LR) and Support Vector Machine Classifier (SVM) from scikit-learn [41]. These models take the flattened upper-triangle connectivity matrix as vector input, instead of using the brain network. (2) General-purposed GNNs: GCN [22], a mean pooling baseline with a graph convolution network as a message-passing layer and Transformer [48], a graph Transformer with mean pooling by taking the connectivity matrix as input. (3) Single-Atlas Models tailored for brain networks: BrainNetCNN [20], the pioneering CNN regressor for connectome data; MG2G [56], a two-stage method with an unsupervised stochastic graph embedding model; and ContrastPool [55], a node clustering pooling using a dual-attention block for domain-specific information capturing. (4) Multi-atlas models: MultiLR, multi-atlas version of LR, concatenate the flatten feature of multiple atlases as input; MultiSVM, multi-atlas version of SVM, similar with MultiLR; MGRL [7], a mean pooling baseline by individual GCN encoder with late fusion; MGT, a multi-atlas version of Transformer with the same fusion mechanism as MGRL; METAFormer [35], a multi-atlas enhanced Transformer with self-supervised pre-training; and LeeNet [27] a multi-atlas GCN approach with early-late fusion.
5.3 Main Results
@@
Table 2: Graph Classification Results (Average Accuracy ± Standard Deviation) over 10-fold-CV. The first and second best results on each dataset are highlighted in bold and underline.
@@
Schaefer100 LR 64.81 ± 3.70 61.97 ± 4.24 56.48 ± 6.76 60.00 ± 20.00
SVM 64.41 ± 5.09 61.52 ± 4.95 63.21 ± 8.62 56.67 ± 17.00
GCN 60.19 ± 2.96 60.40 ± 4.89 54.02 ± 9.06 56.67 ± 17.00
Transformer 59.90 ± 3.77 63.64 ± 2.61 59.33 ± 5.68 60.00 ± 20.00
BrainNetCNN 65.75 ± 3.24 60.48 ± 3.29 57.33 ± 10.32 61.67 ± 13.33
MG2G 64.41 ± 2.16 63.64 ± 5.10 55.45 ± 10.24 61.67 ± 19.79
ContrastPool 65.01 ± 3.84 65.67 ± 6.64 64.00 ± 6.63 61.67 ± 13.02
AAL116 LR 63.80 ± 3.00 64.06 ± 1.80 56.00 ± 7.79 66.67 ± 21.08
SVM 65.72 ± 3.30 63.40 ± 1.90 64.12 ± 5.69 65.00 ± 20.34
GCN 60.10 ± 5.74 61.24 ± 2.47 53.14 ± 8.82 65.00 ± 21.67
Transformer 60.88 ± 4.39 63.27 ± 2.79 61.24 ± 7.22 63.33 ± 24.49
BrainNetCNN 64.58 ± 6.29 62.52 ± 2.91 51.19 ± 9.24 66.67 ± 18.33
MG2G 62.99 ± 4.01 64.41 ± 2.52 59.71 ± 9.11 70.00 ± 19.44
ContrastPool 64.70 ± 3.26 66.33 ± 4.10 63.56 ± 7.90 65.00 ± 20.82
Schaefer100 + AAL116 MultiLR 65.23 ± 5.13 64.99 ± 2.40 55.00 ± 6.25 56.67 ± 24.94
MultiSVM 64.31 ± 5.24 65.21 ± 2.74 63.60 ± 7.66 58.33 ± 17.08
MGRL 61.56 ± 4.90 62.74 ± 3.55 54.55 ± 10.67 68.33 ± 18.93
MGT 63.32 ± 3.90 63.99 ± 4.34 62.14 ± 9.90 65.00 ± 22.91
METAFormer 61.27 ± 4.05 66.52 ± 2.63 54.02 ± 8.81 61.67 ± 25.87
LeeNet 61.28 ± 3.12 64.63 ± 1.34 60.74 ± 4.39 58.33 ± 17.08
AIDFusion (ours) 66.35 ± 3.26 67.57 ± 2.04 66.00 ± 4.71 75.00 ± 13.44
@@
We report the classification accuracy on 4 brain network datasets over 10-fold cross-validation in Table 2. For certain diseases, the effectiveness/informativeness of different atlases is different. On Matai, all the 7 baselines attain better performance with AAL116 than with Schaefer100. On ADNI, 6 out of 7 baselines also perform better with AAL116. In contrast on ABIDE, 5 out of 7 baselines achieve better results with Schaefer100 than with AAL116. It demonstrates the importance of using multi-atlas for brain network analysis instead of relying on one specific atlas. Moreover, it is evident that the multi-atlas baselines with a simple late fusion mechanism (MGRL and MGT) outperform their respective single-atlas models (GCN and Transformer). This highlights the effectiveness of multi-atlas approaches in enhancing the performance of base models. However, conventional ML models (MultiLR and MultiSVM) fail to outperform their single-atlas versions in some cases, possibly due to their inability to effectively utilize multi-atlas features with simple concatenate fusion.
@@
We can also observe that our proposed AIDFusion consistently outperforms not only all single-atlas methods but also state-of-the-art multi-atlas methods across all datasets. Specifically, AIDFusion achieves improvements over all multi-atlas methods on these four datasets by up to 9.76% ((75.00% - 68.33%) / 68.33% = 9.76% on Mātai). Our model gains larger performance improvement on small datasets (PPMI and Mātai) than on large datasets (ABIDE and ADNI), which meets the intuition that information utilization tends to be more critical in applications with smaller sample sizes. Moreover, the results demonstrate that AIDFusion tends to have lower standard deviations compared to other multi-atlas models, indicating the robustness of AIDFusion. This robustness is particularly desirable in medical applications where consistency and reliability are crucial. Apart from accuracy, we also report other evaluation metrics on the ABIDE dataset in Appendix F. The results also reveal the superiority of AIDFusion. Besides Schaefer100 and AAL116, we also conduct experiments with 3 atlases on ADNI by using HO48 [36], which is discussed in Appendix G. AIDFusion is a clear winner under all multi-atlas settings. The influence when using atlases with various resolutions is also discussed in Appendix G.
5.4 Model Interpretation
In neurodegenerative disorder diagnosing, identifying salient ROIs/connections associated with predictions as potential biomarkers is crucial. In this study, we utilize attention scores from the Transformer layer to generate heat maps for brain networks to interpret our model. We visualize these attention maps using the Nilearn toolbox [1]. Figure 2 presents attention maps for two atlases, where higher attention values mean better classification potential for AD (from the ADNI dataset). We utilized 7 networks [58] to assess the connections between our highlighted ROIs and major networks potentially involved with disorders. ROIs from the AAL that do not overlap with these seven networks are excluded from the heat maps. The top 10 ROIs with the highest attention values are displayed in the brain view. As depicted in the attention maps, attention maps of both Schaefer and AAL atlases identify common connections between the visual network (VIS) and the dorsal attention network (DAN), recognized as key connectivities in AD research [4, 2]. Additionally, atlas-specific connections are highlighted. For example, the attention map of Schaefer atlas emphasizes connections within the default mode network (DMN) corresponding with the observations of  Damoiseaux et al. [13]. Findings on the attention map on AAL are consistent with Agosta et al. [2], showing that AD is associated with connectivities in VIS, especially in frontal networks. These findings suggest that AIDFusion effectively captures complementary information from different atlases. We also find some highlighted ROIs that diverge from conventional neuroscientific understanding. For example, connections between VIS and somatomotor network (SMN) have a high attention weight in AIDFusion on Schaefer atlas, which may imply AD is related to the function of defining the targets of actions and providing feedback for visual activation. This insight has not been identified by existing literature. Discussion of ASD (from the ABIDE dataset) is included in Appendix H.
@@ 5.4 Model Interpretation
5.5 Ablation Study
To inspect the effect of the key components in AIDFusion, we conduct experiments by disabling each of them without modifying other settings. The results on ADNI dataset are reported in Table 3. For inter-atlas message-passing (denoted as “IA-MP” in the table), subject-level consistency and population-level consistency, we disable them by simply removing these modules. When disabling “Disen TF”, we replace the disentangle Transformer and the identity embedding with a vanilla Transformer backbone (denoted as “TF” in the table). When disabling all key components (the first row in the table), our model will degenerate to MGT in Table 2. The results demonstrate that AIDFusion with all important modules enabled achieves the best performance. The component that affect the performance most is the population-level consistency. Besides, all variants of the proposed AIDFusion outperform the MGT baseline, demonstrating the effectiveness of our model design. We also conduct a case study of the role of the incompatible nodes in the disentangle Transformer in Appendix I. Further hyperparameter sensitivity analysis is provided in Appendix J.
Table 3: Ablation study on the key components of AIDFusion on ADNI, with the best result bold.
Backbone IA-MP Subject-level Consistency Population-level Consistency acc ± std

### v2

@@ IV-C Subject- and Population-level Consistency
TABLE I: Statistics and class information of brain network datasets used in this work.
Dataset Condition Sex (F/M) Age (mean ± std) Subject# Class#
ABIDE Autism Spectrum Disorder 152/873 16.5 ± 7.4 1025 2
ADNI Alzheimer’s Disease 728/599 74.6 ± 7.9 1326 6
PPMI Parkinson’s Disease 82/127 62.9 ± 9.5 209 4
Matai Mild Traumatic Brain Injury (mTBI) N/A N/A 60 2
A new feature matrix $\hat{\boldsymbol{H}}\in\mathbb{R}^{n^{\prime}\times d}$ is generated by the cluster assignment matrix $\boldsymbol{S}$ and the embedded node feature matrix $\boldsymbol{Z}$ as $\hat{\boldsymbol{H}}=\boldsymbol{S}^{\mathsf{T}}\boldsymbol{Z}$. This coarsening process aims to generate super-nodes to get higher-level node representations. To avoid GNN treating each ROI and each node cluster equally, we adopt an entropy regularization to the assignment matrices of each atlas:
$\small\mathcal{L}_{E}=\frac{1}{n^{\prime}}\sum^{n^{\prime}}_{{\color[rgb]{0,0,0}j}=1}\left(\operatorname{entropy}(\boldsymbol{S}^{a}[{\color[rgb]{0,0,0}j},:])+\operatorname{entropy}(\boldsymbol{S}^{b}[{\color[rgb]{0,0,0}j},:])\right),$ (9)
$\small\operatorname{entropy}(\boldsymbol{p})=-\sum_{j=1}^{n^{\prime}}\boldsymbol{p}_{j}\log(\boldsymbol{p}_{j}).$ (10)
We elaborate on the module of subject-level consistency in Figure 1d. Through two DiffPool layers, we align the number of nodes and produce high-quality representations for each atlas. Then we are able to apply a contrastive loss to them by considering representations from the same subject as positive pairs $\mathcal{P}^{pos}=\{(\hat{\boldsymbol{H}}^{a}_{i},\hat{\boldsymbol{H}}^{b}_{i})\}:i=1,\dots,bz\}$ and representations from different subjects as negative pairs $\mathcal{P}^{neg}=\{(\hat{\boldsymbol{H}}^{a}_{i},\hat{\boldsymbol{H}}^{b}_{\neg i}):i=1,\dots,bz\}$:
@@
Population-level Consistency. The readout function $\boldsymbol{m}=\operatorname{READOUT}\left(\boldsymbol{H}\right)$ is an essential component of learning the graph-level representations $\boldsymbol{m}\in\mathbb{R}^{d}$ for brain network analysis (e.g., classification), which maps a set of learned node-level embeddings to a graph-level embedding. To further constrain the consistency for graph representations across different atlases, we introduce a mean squared error (MSE) loss on the population level. Different from some existing works, such as FTMMR [56], that directly enforce the alignment of different views on graph level, we focus on maintaining the relative similarity structure among subjects across atlases. As shown in Figure 1e, a population graph $\boldsymbol{G}$ is constructed by computing the similarity of each two subjects’ graph representations in the same atlas. The intuition here is we aim to maintain the relationship of subjects across atlases, instead of directly enforcing graph representations of two atlases to be the same. This constraint does not rely on the target label, as we only encourage the consistency of (dis)similarity across atlases rather than enforcing the similarity within each labeled group. Such loss is formulated as follows:
@@
Total Loss. The model is supervised by a widely-used cross-entropy loss $\mathcal{L}_{cls}$ [57] for brain network classification. The total loss is computed by:
@@
V Experimental Results
V-A Brain Network Datasets
We use four brain network datasets released by Xu et al. [42] to study various disorders: ABIDE [58] for Autism Spectrum Disorder (ASD), ADNI [59] for Alzheimer’s Disease (AD), PPMI [60] for Parkinson’s Disease (PD), and Mātai for mild traumatic brain injury (mTBI) [42]. A summary of the brain network datasets is provided in Table I. The atlases used in this study are Schaefer [10] and AAL [8], with 100 and 116 ROIs, respectively. We choose these two atlases because they are among the most widely used and well-established brain parcellation schemes in the neuroimaging community. Importantly, they represent two complementary perspectives: the AAL atlas is based on anatomical landmarks, while the Schaefer atlas is derived from functional connectivity patterns.
ABIDE The ABIDE initiative aggregates functional brain imaging data from multiple research labs worldwide to support ASD research. The age range of participants spans from 5 to 64 years. The dataset classifies subjects into two groups: typical controls (TC) and individuals diagnosed with ASD.
ADNI The raw images used in this study were obtained from the ADNI database (adni.loni.usc.edu). Launched in 2003 as a public-private partnership led by Principal Investigator Michael W. Weiner, MD, the primary goal of ADNI is to test whether serial MRI, PET, other biological markers, and clinical assessments can be combined to measure the progression of mild cognitive impairment (MCI) and early Alzheimer’s Disease (AD). For up-to-date information, visit www.adni-info.org. We include subjects from six distinct stages of AD: cognitively normal (CN), significant memory concern (SMC), MCI, early MCI (EMCI), late MCI (LMCI), and AD.
PPMI The Parkinson’s Progression Markers Initiative (PPMI) is a comprehensive study aimed at identifying biological markers associated with Parkinson’s risk, onset, and progression. The PPMI dataset includes multimodal and multi-site MRI images from subjects in four distinct groups: normal controls (NC), scans without evidence of dopaminergic deficit (SWEDD), prodromal Parkinson’s disease, and diagnosed PD.
Mātai The Mātai study is a longitudinal, single-site, single-scanner investigation designed to detect subtle brain changes due to a season of contact sports. This dataset consists of brain network data collected from the Gisborne-Tairāwhiti region of New Zealand, with 35 contact sport players imaged pre-season (N=35) and post-season (N=25). Subtle brain changes confirmed by diffusion imaging studies were observed as a result of playing contact sports.
V-B Baseline Models
@@
We use 8 single-atlas methods and 6 multi-atlas methods as baselines to evaluate our proposed AIDFusion, including: (1) Conventional machine learning (ML) models: Logistic Regression (LR) and Support Vector Machine Classifier (SVM) from scikit-learn [61]. These models take the flattened upper-triangle connectivity matrix as vector input, instead of using the brain network. (2) General-purposed GNNs: GCN [53] and Transformer [32]. (3) Single-Atlas Models tailored for brain networks: BrainNetCNN [5], MG2G [29], ContrastPool [31] and BNT [36]. (4) Multi-atlas models: MultiLR (multi-atlas version of LR, concatenate the flatten feature of multiple atlases as input), MultiSVM (multi-atlas version of SVM, similar with MultiLR); MGRL [12]; MGT (a multi-atlas version of Transformer with the same fusion mechanism as MGRL), METAFormer [13] and LeeNet [14].
The experimental settings are based on those outlined in Xu et al. [31]. For each dataset, we use an 8:1:1 split for training, validation, and testing, respectively. Each model is evaluated across 5 random seeds using 10-fold cross-validation (CV), with the average accuracy reported for the main results. We apply an early stopping criterion: the learning rate is halved if there is no improvement in validation loss over 25 epochs, and training stops once the learning rate reaches a pre-defined minimum threshold. All experiments were conducted on a server equipped with an AMD Ryzen Threadripper PRO 5995WX (64 cores) and an NVIDIA GeForce RTX 4090.
All the baselines used in this paper are implemented by ourselves. In AIDFusion, we adopt a mean pooling layer as the readout function and a two-layer MLP with ReLU as the prediction head. The number of incompatible nodes $r$ is set to 4. The number of clusters $n^{\prime}$ for the cluster assignment matrix $\boldsymbol{S}$ is set to half of the average number of nodes for all atlases, e.g., for Schaefer100 and AAL116, $n^{\prime}=(100+116)/2/2=54$. The temperature hyper-parameter $\tau$ in Eq. (11) is set to 0.75. The hidden dimensions of all layers are set to 100. The batch size of each dataset is set to 10% of the subject number in the dataset. We tuned the other hyperparameters on validation set including $k$ in kkNN for the inter-atlas message-passing, trade-off hyperparameters $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ for loss function Eq. (13), the initial learning rate init_lrinit\_lr, and the minimum learning rate min_lrmin\_lr used for early stop. To be specific, we search $k$ from {3, 5, 10}, $\lambda_{1}$ from {1e-2, 1e-1, 1e0, 1e1, 1e2}, $\lambda_{2}$ from {1e-2, 1e-1, 1e0, 1e1, 1e2}, $\lambda_{3}$ from {1e-6, 1e-5, 1e-4, 1e-3, 1e-2}, $\lambda_{4}$ from {1e-1, 1e0, 1e1}, inir_lrinir\_lr from {5e-5, 8e-5, 1e-4, 2e-4, 7e-4} and inir_lrinir\_lr from {6e-5, 1e-5, 5e-6, 1e-6}. The optimized hyperparameters for AIDFusion are reported in Table II.
TABLE II: The optimized hyperparameters for AIDFusion.
ABIDE ADNI PPMI Mātai
$k$ 10 5 5 5
$\lambda_{1}$ 1e1 1e1 1e1 1e-2
$\lambda_{2}$ 1e-1 1e1 1e1 1e0
$\lambda_{3}$ 1e-2 1e-5 1e-4 1e-3
$\lambda_{4}$ 1e0 1e0 1e0 1e0
init_lrinit\_lr 1e-4 8e-5 1e-4 7e-4
min_lrmin\_lr 6e-5 6e-5 6e-5 6e-5
For machine learning baselines (LR, SVM), we use the official code of Xu et al. [42] and implement the multi-atlas version on the top of it. We follow the same parameter-tuning scheme as they do. The full list of parameters, including those we used for the grid search, is provided below. The parameter tuning for MultiLR and MultiSVM are the same with LR and SVM, respectively.
@@
TABLE III: Brain Network Classification Results (Average Accuracy ± Standard Deviation) over 5 runs. The first and second best results on each dataset are highlighted in bold and underline. The p-values of one-sided paired t-tests comparing our AIDFusion with the best baselines on the four datasets are 0.0483, 0.0040, 0.0141, and 0.0068, respectively.
@@
Schaefer100 LR 64.81 ± 0.00 61.97 ± 0.00 56.48 ± 0.00 60.00 ± 0.00
SVM 64.41 ± 0.00 61.52 ± 0.00 63.21 ± 0.00 56.67 ± 0.00
GCN 60.16 ± 1.76 60.62 ± 0.42 52.60 ± 1.58 60.33 ± 3.80
Transformer 60.34 ± 1.75 63.91 ± 0.82 58.67 ± 1.94 60.00 ± 3.73
BrainNetCNN 64.93 ± 0.79 63.42 ± 1.21 53.53 ± 2.26 60.33 ± 7.94
MG2G 63.51 ± 0.30 62.98 ± 0.39 56.46 ± 1.45 61.67 ± 2.36
ContrastPool 65.40 ± 1.01 65.13 ± 0.96 63.79 ± 1.88 62.00 ± 4.77
BNT 60.56 ± 1.19 66.14 ± 0.65 57.31 ± 1.37 60.67 ± 2.79
AAL116 LR 63.80 ± 0.00 64.06 ± 0.00 56.00 ± 0.00 66.67 ± 0.00
SVM 65.72 ± 0.00 63.40 ± 0.00 64.12 ± 0.00 65.00 ± 0.00
GCN 59.50 ± 0.62 61.02 ± 0.75 51.20 ± 1.63 56.67 ± 4.25
Transformer 60.05 ± 0.78 63.52 ± 0.47 60.87 ± 1.88 62.67 ± 4.18
BrainNetCNN 63.94 ± 1.00 63.09 ± 0.90 50.79 ± 1.04 68.33 ± 2.64
MG2G 63.32 ± 1.40 63.34 ± 0.70 58.76 ± 1.72 69.00 ± 3.03
ContrastPool 64.92 ± 0.90 66.18 ± 1.13 63.22 ± 1.35 64.67 ± 3.21
BNT 59.65 ± 1.44 60.73 ± 0.87 55.63 ± 1.76 63.00 ± 4.63
Schaefer100
AAL116
MultiLR 65.23 ± 0.00 64.99 ± 0.00 55.00 ± 0.00 56.67 ± 0.00
MultiSVM 64.31 ± 0.00 65.21 ± 0.00 63.60 ± 0.00 58.33 ± 0.00
MGRL 60.14 ± 0.94 62.65 ± 0.90 54.37 ± 1.37 68.66 ± 3.61
MGT 63.71 ± 1.73 63.36 ± 0.49 62.62 ± 1.55 65.33 ± 3.21
METAFormer 61.97 ± 0.87 66.76 ± 0.28 54.82 ± 1.91 67.64 ± 6.70
LeeNet 61.33 ± 0.54 64.60 ± 0.17 60.83 ± 0.21 58.33 ± 0.0
AIDFusion (ours) 66.51 ± 1.45 67.78 ± 0.35 66.04 ± 1.28 73.67 ± 1.39
TABLE IV: Results of more evaluation metrics (precision, recall, micro-F1, and ROC-AUC) on ABIDE dataset over 5 runs. The best result is highlighted in bold. The p-values of one-sided paired t-tests for all metrics except precision are under 0.05.
Precision Recall micro-F1 ROC-AUC
MGRL 59.81 ± 3.12 59.79 ± 3.05 59.59 ± 2.98 61.36 ± 2.01
MGT 60.89 ± 3.21 67.75 ± 5.68 64.08 ± 1.77 61.47 ± 1.52
METAFormer 60.06 ± 0.98 61.31 ± 0.61 60.52 ± 0.67 61.94 ± 0.83
LeeNet 64.50 ± 0.86 43.53 ± 0.99 51.58 ± 0.99 60.52 ± 0.55
AIDFusion 62.31 ± 3.02 73.71 ± 6.03 67.69 ± 1.67 66.57 ± 1.15
TABLE V: Results of more atlases on ABIDE dataset over 10-fold-CV in a single run. The best results for each atlas setting are highlighted in bold. The p-value of one-sided paired t-test for the 3-atlas setting is under 0.1, while p-values for the other 2-atlas settings are under 0.05.
Atlas model acc ± std
Schaefer100 AAL116 BASC122
MGRL 61.56 ± 4.90
MGT 63.32 ± 3.90
METAFormer 61.27 ± 4.05
LeeNet 61.28 ± 3.12
AIDFusion 66.35 ± 3.26
MGRL 60.80 ± 5.12
MGT 58.49 ± 5.64
METAFormer 62.92 ± 5.79
LeeNet 60.01 ± 4.00
AIDFusion 65.97 ± 3.60
MGRL 63.14 ± 4.17
MGT 59.90 ± 3.46
METAFormer 62.53 ± 3.78
LeeNet 59.12 ± 4.20
AIDFusion 64.79 ± 2.80
MGRL 63.62 ± 4.57
MGT 62.84 ± 3.85
METAFormer 65.80 ± 5.61
LeeNet 60.19 ± 3.77
AIDFusion 66.65 ± 4.14
V-C Main Results
@@
The classification accuracy on 4 brain network datasets over 5 runs is reported in Table III. For certain diseases, the effectiveness/informativeness of different atlases is different. On Matai, 6 out of 7 baselines attain better performance with AAL116 than with Schaefer100. On ADNI, 4 out of 7 baselines also perform better with AAL116. In contrast on ABIDE, 6 out of 7 baselines achieve better results with Schaefer100 than with AAL116. It demonstrates the importance of using multi-atlas for brain network analysis instead of relying on one specific atlas. Moreover, it is evident that the multi-atlas baselines with a simple late fusion mechanism (MGRL and MGT) outperform their respective single-atlas models (GCN and Transformer) in the most cases. This highlights the effectiveness of multi-atlas approaches in enhancing the performance of base models. However, conventional ML models (MultiLR and MultiSVM) fail to outperform their single-atlas versions in more cases, possibly due to their inability to effectively utilize multi-atlas features with simple concatenate fusion.
@@
We can also observe that our proposed AIDFusion consistently outperforms not only all single-atlas methods but also the state-of-the-art multi-atlas methods across all these datasets. Specifically, AIDFusion achieves improvements over all multi-atlas methods on these four datasets by up to 6.77% ((73.67% - 69.00%) / 69.00% = 6.77% on Mātai). Our model gains more improvement on small datasets (PPMI and Mātai) than on large datasets (ABIDE and ADNI), which meets the intuition that information utilization tends to be more critical in applications with smaller sample sizes.
In addition to accuracy, we also report other evaluation metrics, including precision, recall, micro-F1, and ROC-AUC, for all the multi-atlas deep models on the ABIDE dataset. As displayed in Table IV, the proposed AIDFusion performs the best across all these evaluation metrics except for precision. We observe that compared to other baselines, our AIDFusion can significantly improve recall without compromising precision. Moreover, in medical diagnostics, it is crucial to ensure that all individuals with a certain condition are correctly identified, even if it leads to some false positives. Missing a true positive (failing to diagnose a disease) can have severe consequences, while false positives can be further examined or retested. Therefore, models with higher recall rates, such as our AIDFusion, are more suitable for real-life medical auxiliary diagnosis.
@@
Besides Schaefer100 and AAL116, we also conduct experiments with 3 atlases on ABIDE by using BASC122 [62]). When extending AIDFusion to scenarios involving three atlases, each component of our framework generalizes naturally from the two-atlas case. Specifically, inter-atlas message-passing is applied independently to each pair of atlases. For three atlases (e.g., A, B, and C), we construct inter-atlas edges and perform message-passing for the pairs (A, B), (A, C), and (B, C). Similarly, the subject-level and population-level consistency losses are computed for each atlas pair and then averaged across all combinations to form the final loss terms. Results shown in Table V demonstrate that for five multi-atlas methods, three of them (MGRL, METAFormer, and AIDFusion) achieve better performance with three atlases compared to any two-atlas combinations. Importantly, our proposed AIDFusion still outperforms all baselines in these settings. This reinforces our claim that AIDFusion effectively integrates multi-atlas information while remaining robust to the inclusion of additional atlases. Besides, Table V also shows METAFormer exhibiting a large performance gain when incorporating three atlases. However, this gain is primarily due to its relatively low performance in the two-atlas setting, rather than an indication of superior scalability. In fact, even with three atlases, METAFormer still underperforms compared to AIDFusion with only two atlases. This further highlights the effectiveness of our model in extracting comprehensive subject-level representations without requiring excessive atlas redundancy.
V-D Ablation Study
To inspect the effect of the key components in AIDFusion, we conduct experiments by disabling each of them without modifying other settings. The results on ADNI dataset are reported in Table VI. For inter-atlas message-passing (denoted as “IA-MP” in the table), subject-level consistency (SC) and population-level consistency (PC), we disable them by simply removing these modules. When disabling “Disen TF”, we replace the disentangle Transformer and the identity embedding with a vanilla Transformer backbone (denoted as “TF” in the table). We further conduct experiments by disabling the identity embedding (shown as “Disen TF w/o ID”), and the orthogonal loss (shown as “Disen TF w/o $L_{orth}$”) described in Section IV-A. When disabling all key components (the first row in the table), our model will degenerate to MGT in Table III. The results demonstrate that AIDFusion with all important modules enabled achieves the best performance. The component that affects the performance most is the population-level consistency. Besides, all variants of the proposed AIDFusion outperform the MGT baseline, demonstrating the effectiveness of our model design. This table highlights not only the additive value of each module but also the superiority of our framework compared to existing multi-atlas models, even when some components are removed.
TABLE VI: Ablation study on the key components of AIDFusion on ADNI over 10-fold-CV of a single run, with the best result bold.
Backbone IA-MP SC PC acc ± std